# Mingkai Liao

### v1

@@ Abstract
Opinion polarization in online social networks poses a critical challenge to social cohesion and democratic processes. Recent work formulates polarization moderation as algorithmic intervention problems under opinion dynamics models, most notably the Friedkin–Johnsen (FJ) model. However, existing approaches are largely tailored to specific linear settings and rely on closed-form steady-state analysis, limiting their scalability, adaptability, and applicability to cost-aware, nonlinear, or topology-altering interventions.
In this paper, we propose PACIFIER, a graph reinforcement learning framework for sequential polarization moderation via network interventions. PACIFIER reformulates the canonical ModerateInternal (MI) and ModerateExpressed (ME) problems as sequential decision-making tasks, enabling adaptive policies without repeated recomputation of steady-state solutions. The framework is objective-agnostic and naturally extends to a broad class of FJ-consistent settings, including budget-aware interventions, continuous internal opinions, biased-assimilation dynamics, and node removal.
We introduce a temporal-aware node marking mechanism to encode intervention history under fixed topology, together with polarization-aware global features that correlate with polarization regimes while avoiding expensive steady-state recomputation during planning. Extensive experiments on 15 real-world Twitter follow and retweet networks (up to 155,599 nodes) demonstrate strong and consistent performance. In expressive-opinion moderation tasks (ME and ME-cost), PACIFIER-RL achieves a 100% win rate against the strongest non-learning baselines, yielding average improvements of 15%–40% in AUC. Under cost-aware MI settings, both RL and greedy variants consistently dominate all baselines with nearly 40% average gains. Even under nonlinear biased-assimilation dynamics and topology-changing node removal interventions, PACIFIER-RL maintains robust superiority, while in linear MI settings it performs competitively with near-optimal analytical methods.
These results establish graph reinforcement learning as a unified, scalable, and robust paradigm for pacing and moderating opinion polarization across diverse dynamical regimes and intervention mechanisms.
@@ 1 Introduction
The primary contribution of this work is the introduction of PACIFIER, a graph reinforcement learning framework for sequential polarization moderation. We reformulate the MI and ME problems as sequential decision-making tasks over networked opinion dynamics, enabling the learning of adaptive intervention policies without relying on repeated steady-state recomputation or handcrafted heuristics. Unlike prior analytical approaches, PACIFIER is objective-agnostic: the agent optimizes any polarization-related objective specified through its reward function. Within this unified paradigm, we demonstrate that PACIFIER naturally extends beyond the original linear MI/ME settings to several practically relevant regimes, including cost-aware intervention, continuous-valued internal opinions, biased-assimilation nonlinear dynamics, and topology-altering node removal.
@@ 1.1 Contributions
  • Generalization across intervention paradigms and FJ-consistent dynamics. Within a single objective-agnostic learning framework, PACIFIER extends beyond the original linear MI/ME settings to multiple practically relevant regimes, including cost-aware interventions, continuous-valued internal opinions, homogeneous nonlinear biased-assimilation dynamics, and topology-altering node removal. This establishes a unified moderation paradigm that remains robust under structurally consistent extensions of the Friedkin–Johnsen model.
@@ 2.4 Research Gap and Our Contribution
PACIFIER bridges this gap by formulating ModerateInternal and ModerateExpressed—and their FJ-consistent extensions—as sequential decision problems over networked opinion dynamics. Our framework is objective-agnostic and accommodates multiple intervention paradigms, including cost-aware moderation, biased-assimilation dynamics, and node removal, while preserving the underlying influence topology when appropriate. To address topology-preserving aliasing, we introduce a temporal-aware node marking mechanism, and to enhance value estimation, we incorporate polarization-aware global features.
@@ 3.2 Polarization Index
3.3 Canonical Moderation Problems
We now formalize the two key intervention problems for reducing polarization. Throughout, we assume a budget $k$, and the goal is to select $k$ individuals to moderate, i.e., to set the targeted opinion attribute to a neutral state (0), such that the resulting polarization index is minimized.
Base vs. continuous opinion settings.
We distinguish two closely related opinion settings. (i) Base setting (Matakos et al.). The internal opinions are binary and polarized, i.e., $s_{i}\in\{-1,+1\}$, where 1-1 and +1+1 represent two opposing camps and 0 denotes neutrality after moderation. (ii) Continuous setting (our extension). The internal opinions are continuous, $s_{i}\in[-1,1]$, which allows heterogeneous opinion intensities. The two opposing camps correspond to negative vs. positive opinions, i.e., $s_{i}\in(-1,0)$ and $s_{i}\in(0,1)$, with $s_{i}=0$ representing the neutral stance. In both settings, the moderation action sets the targeted attribute to 0 for selected individuals; the continuous setting generalizes the base setting by allowing magnitudes between the extremes.
@@ 3.3.2 The ModerateExpressed Problem
3.4 Our Setting: One-Shot Planning with ANP Evaluation
The canonical MI/ME formulations above specify what is being intervened on, but do not specify what information a planning algorithm may use while constructing the intervention trajectory. Here we study a setting with two coupled components: a one-shot planning constraint and an accumulated normalized polarization (ANP) evaluation metric.
(i) One-shot planning constraint (plan once from the initial instance).
@@ 3.4 Our Setting: One-Shot Planning with ANP Evaluation
(ii) Accumulated normalized polarization (ANP) evaluation metric (area-under-curve).
@@
so the full sequence $(v_{1},\dots,v_{k})$ traces a polarization trajectory from $x_{0}=0$ (no intervention) to $x_{k}=1$ (all $k$ interventions executed). We define the normalized polarization at step $t$ as
$\widehat{\pi}\!\left(\mathbf{z}^{(t)}\right)=\frac{\pi\!\left(\mathbf{z}^{(t)}\right)}{n},$ (5)
where $\pi(\mathbf{z})=\|\mathbf{z}\|_{2}^{2}/n$ is the polarization index in (4). Then the accumulated normalized polarization (ANP) score is defined as the discrete area-under-curve proxy:
$\mathcal{R}_{\mathrm{ANP}}(v_{1},\dots,v_{k})=\frac{1}{k}\sum_{t=0}^{k}\widehat{\pi}\!\left(\mathbf{z}^{(t)}\right).$ (6)
Intuitively, $\mathcal{R}_{\mathrm{ANP}}$ approximates the area under the polarization curve $\widehat{\pi}(\mathbf{z}^{(t)})$ over the normalized horizon $x\in[0,1]$. Lower values indicate that polarization is reduced earlier (small values for small $t$) and persistently (small values throughout the trajectory), rather than only near the end.
Implication.
Our setting is sequential (since the output is an ordered trajectory and every prefix contributes to evaluation), but the planner is one-shot: it commits to the entire sequence based only on initial instance information, and the sequence is then directly evaluated under the ANP (area-under-curve) criterion emphasizing when polarization reduction occurs.
@@ 3.5 Computational Complexity
3.6 Extensibility Beyond the Linear FJ Model
The formulations above are presented under the linear Friedkin–Johnsen (FJ) model for clarity. However, a central motivation of this work is to develop a moderation framework that is not tied to a specific analytical form of the opinion dynamics. In this subsection, we clarify the class of extensions considered in this paper and illustrate how PACIFIER naturally generalizes beyond the linear FJ setting.
FJ-consistent extensions.
We focus on extensions that preserve the core social influence structure of the FJ model—namely, influence propagation along the edges of the social network and the distinction between internal and expressed opinions—while relaxing either the network topology or the opinion update rule. Importantly, these extensions may break linearity and analytical tractability, but remain compatible with PACIFIER because the framework does not rely on closed-form solutions.
Topology-altering interventions: node removal.
As a first class of extensions, we consider Topology-altering interventions that explicitly modify the network topology. In particular, removing a node eliminates both its incoming and outgoing influence, resulting in a dynamically changing graph structure. This setting is analogous to classical network dismantling problems and serves as a bridge between opinion moderation and structural network intervention. Unlike MI and ME, where the topology remains fixed, node removal alters the Laplacian matrix itself and invalidates precomputed influence matrices such as $Q=(L+I)^{-1}$. PACIFIER naturally accommodates this setting by operating directly on the evolving graph state rather than relying on static influence operators.
Dynamics-altering interventions: biased assimilation.
As a second class of extensions, we consider nonlinear variants that retain the original topology but modify the opinion update rule. A representative example is the biased assimilation model proposed by [3], which captures the tendency of individuals to overweight information that aligns with their current beliefs. The update rule for an individual $i$ is given by
$x_{i}(t+1)=\frac{w_{ii}x_{i}(t)+(x_{i}(t))^{b}s_{i}(t)}{w_{ii}+(x_{i}(t))^{b}s_{i}(t)+(1-x_{i}(t))^{b}(d_{i}-s_{i}(t))},$ (7)
where $b\geq 0$ controls the strength of confirmation bias. For $b\geq 1$, this model generates strong polarization effects but generally lacks a closed-form steady-state solution.
Implications for PACIFIER.
Both extensions highlight a key limitation of model-specific moderation algorithms: once linearity is lost or interventions alter the topology, approaches that rely on repeated matrix inversions (such as BOMP) or analytical steady-state expressions (such as ExtremeExpressed) become infeasible. In contrast, PACIFIER treats opinion dynamics as a black-box environment and evaluates interventions solely through their observed impact on polarization. As a result, the same learning-based framework can be applied without modification to linear and nonlinear dynamics, as well as to topology-preserving and topology-altering interventions. These extensions therefore serve not as separate problem formulations, but as evidence of the generality and flexibility of the proposed approach.
@@ 4 PACIFIER Framework
PACIFIER is a graph-based learning framework for sequential polarization moderation on networks. It is designed to learn intervention policies from simulated interactions and deploy them efficiently on large real-world graphs. The framework follows a two-phase workflow: (i) offline training on synthetic graph–opinion instances, and (ii) feed-forward deployment on target networks without iterative recomputation during planning.
@@
PACIFIER supports two closely related variants within the same architecture: PACIFIER-RL, which learns a long-horizon intervention policy via bootstrapped reinforcement learning, and PACIFIER-Greedy, a myopic variant that learns to approximate immediate polarization reduction without bootstrapping. Both variants share the same encoder–decoder representation and differ only in their training objectives.
@@
Design principles. Polarization moderation under MI/ME-type interventions differs fundamentally from topology-altering structural control. In many practical settings, the graph topology remains fixed, while polarization evolves through opinion dynamics and intervention history. This shift motivates three core design principles underlying PACIFIER:
  • Structured episode generation (two-echo-chamber initialization). Each training episode is initialized from a synthetic graph with two communities (echo chambers) and an aligned opinion assignment. Nodes are partitioned into $V^{+}$ and $V^{-}$, dense within-group connections are generated with controllable sparse cross-group exposure, and internal opinions are initialized according to community membership (e.g., $s_{i}^{(0)}=+1$ for $i\in V^{+}$ and $s_{i}^{(0)}=-1$ for $i\in V^{-}$). Varying the inter-group edge ratio yields controllable polarization regimes and exposes the policy to diverse yet structured states.
  • Trajectory-aware reward design (optionally cost-weighted). Rather than optimizing only a terminal objective, PACIFIER assigns stepwise rewards that encourage reducing polarization early along the intervention sequence. Node-specific intervention costs can be incorporated naturally, allowing the framework to balance polarization reduction against heterogeneous moderation effort.
  • History-aware and polarization-aware representations. Because the topology may remain unchanged across decision steps, intervention history cannot be inferred from structure alone. PACIFIER therefore explicitly encodes intervention history at the node level to prevent history-induced state aliasing. In addition, it augments learned graph embeddings with polarization-related auxiliary features that capture opinion–structure interaction patterns, providing value-relevant global signals for stable policy learning.
These components jointly enable PACIFIER to learn scalable intervention strategies that generalize from small synthetic graphs to large real-world networks while maintaining lightweight inference at deployment time.
4.1 Framework
We follow the inductive graph-RL paradigm of FINDER and adopt the same two-phase workflow: (i) offline training on synthetic instances, and (ii) online (deployment-time) application on a target real network, where inference is a lightweight feed-forward scoring of all candidate nodes.
We instantiate two closely related variants within the same architecture: PACIFIER-RL, which adopts multi-step Q-learning with bootstrapping, and PACIFIER-Greedy, a simplified myopic variant that learns only from immediate rewards without bootstrapping.
MDP formulation.
State.
At step $t$, the state is
$s_{t}=(G,X_{t},\mathbf{u}_{t},\mathbf{m}_{t}),$
where $G=(V,E)$ is the (fixed) network, $X_{t}\in\mathbb{R}^{n\times d_{x}}$ is the node-feature matrix, $\mathbf{u}_{t}\in\mathbb{R}^{d_{u}}$ is a graph-level auxiliary feature vector, and $\mathbf{m}_{t}\in\{0,1\}^{n}$ is an action-feasibility mask. Let $S_{t}\subseteq V$ denote the set of already intervened nodes. Then $\mathbf{m}_{t}(v)=0$ for $v\in S_{t}$ and $\mathbf{m}_{t}(v)=1$ otherwise.
Under MI/ME-style moderation, the topology of $G$ remains fixed. State evolution is therefore driven by opinion updates and intervention history, which is explicitly encoded in $X_{t}$ (Sec. 4.3).
Action.
An action selects one feasible node:
$a_{t}\in\mathcal{A}_{t}=\{v\in V\mid\mathbf{m}_{t}(v)=1\}.$
Transition.
Reward (trajectory-aware signal).
$r_{t}=-\,\frac{\pi(\mathbf{z}^{(t)})}{C}\cdot c(a_{t}).$
For unweighted settings, $c(v)\equiv 1$.
Policy parameterization via encode–decode.
Offline training.
PACIFIER-RL. During offline training, PACIFIER-RL interacts with sampled synthetic instances from $\mathcal{D}$ to generate trajectories $\{(s_{t},a_{t},r_{t},s_{t+1})\}$ over a fixed horizon (budget) $k$ (or until no feasible action remains). At each step, the agent observes the state representation $(G,X_{t},\mathbf{u}_{t},\mathbf{m}_{t})$ and selects one feasible node via an $\epsilon$-greedy strategy: with probability $1-\epsilon$ it chooses the feasible node with the largest predicted $Q_{\theta}(s_{t},v)$, and otherwise samples a feasible node uniformly at random. The exploration rate $\epsilon$ is annealed from a high initial value to a small final value over training episodes.
$R_{t:t+n}=\sum_{i=0}^{n-1}\gamma^{i}r_{t+i}.$
$y_{t}=R_{t:t+n}+\gamma^{n}\max_{v}Q_{\theta^{-}}(s_{t+n},v),$
$y_{t}=r_{t}.$
Parameters are optimized by minimizing
Online application (deployment).
After offline training, both variants are deployed on a target real-world graph in a purely feed-forward manner. At each decision step $t$, we construct the current state representation $(G,X_{t},\mathbf{u}_{t},\mathbf{m}_{t})$ and evaluate $Q_{\theta}(s_{t},v)$ for all nodes in parallel. Feasibility is enforced by the mask $\mathbf{m}_{t}$: nodes with $\mathbf{m}_{t}(v)=0$ (already intervened) are assigned a large negative value and are never selected. The deployed policy is greedy with respect to $Q_{\theta}$:

### v2

@@ Abstract
Opinion polarization moderation has been studied predominantly as an analytical optimization problem under the Friedkin–Johnsen (FJ) model, where intervention algorithms are tightly coupled to linear steady-state analysis and model-specific derivations. While effective in narrowly structured settings, such methods scale poorly and do not naturally extend to richer intervention regimes. This raises a central question: can polarization moderation be treated not merely as an analytical domain, but as a graph-based sequential planning problem?
We answer this question by proposing PACIFIER, to our knowledge the first unified graph-learning framework, and in particular the first graph reinforcement learning framework, for opinion polarization moderation. PACIFIER reformulates the canonical ModerateInternal (MI) and ModerateExpressed (ME) problems as sequential decision-making tasks on graphs, replacing repeated analytical recomputation with learned intervention policies. The framework is instantiated with two complementary variants: PACIFIER-RL, which learns long-horizon intervention policies through reinforcement learning, and PACIFIER-Greedy, which learns efficient myopic action ranking. Within the same architecture, PACIFIER extends naturally beyond the canonical setting to cost-aware moderation, continuous-valued internal opinions, and topology-altering node removal.
To address the distinctive challenges of topology-preserving moderation, we introduce a temporal-aware node marking mechanism that makes intervention history explicit to the encoder, together with polarization-aware global features that summarize opinion–structure interaction patterns. These components provide a more history-aware and value-relevant state representation while avoiding expensive steady-state recomputation during planning. Under an inductive learning paradigm, PACIFIER is trained only on small synthetic two-echo-chamber graphs with fewer than 50 nodes, yet generalizes effectively to real-world Twitter follow and retweet networks with up to 155,599 nodes.
Experiments on 15 real-world polarized networks reveal a clear regime-dependent picture. In analytically structured MI settings, PACIFIER remains competitive with strong analytical solvers and consistently emerges as the strongest scalable non-analytical alternative. In contrast, in ME, continuous-ME, and cost-ME, PACIFIER achieves strong and highly consistent superiority over non-PACIFIER baselines. Most importantly, PACIFIER-RL becomes decisively superior when long-horizon planning truly matters, most notably in cost-ME and topology-altering node_removal, where reasoning over future structural consequences is crucial.
Overall, PACIFIER shifts opinion polarization moderation from a domain dominated by model-specific analytical optimization toward a unified graph-learning and graph-reinforcement-learning paradigm.
@@ 1 Introduction
The primary contribution of this work is the introduction of PACIFIER, a graph reinforcement learning framework for sequential polarization moderation. We reformulate the MI and ME problems as sequential decision-making tasks over networked opinion dynamics, enabling the learning of adaptive intervention policies without relying on repeated steady-state recomputation or handcrafted heuristics. Unlike prior analytical approaches, PACIFIER is objective-agnostic: the agent optimizes any polarization-related objective specified through its reward function. Within this unified paradigm, we demonstrate that PACIFIER naturally extends beyond the original linear MI/ME settings to several practically relevant regimes, including cost-aware intervention, continuous-valued internal opinions, and topology-altering node removal.
@@ 1.1 Contributions
  • Generalization across intervention paradigms and FJ-consistent dynamics. Within a single objective-agnostic learning framework, PACIFIER extends beyond the original linear MI/ME settings to multiple practically relevant regimes, including cost-aware interventions, continuous-valued internal opinions, and topology-altering node removal. This establishes a unified moderation paradigm that remains robust under structurally consistent extensions of the Friedkin–Johnsen model.
@@ 2.4 Research Gap and Our Contribution
PACIFIER bridges this gap by formulating ModerateInternal and ModerateExpressed—and their FJ-consistent extensions—as sequential decision problems over networked opinion dynamics. Our framework is objective-agnostic and accommodates multiple intervention paradigms, including cost-aware moderation, and node removal, while preserving the underlying influence topology when appropriate. To address topology-preserving aliasing, we introduce a temporal-aware node marking mechanism, and to enhance value estimation, we incorporate polarization-aware global features.
@@ 3.2 Polarization Index
This quantity is already normalized by the number of nodes $n$, so that polarization values are comparable across graphs of different sizes.
3.3 Topology-Preserving Interventions
Before introducing the canonical MI and ME problems, we formalize the class of interventions to which they belong.
Definition 3.2 (Topology-preserving intervention).
An intervention process is topology-preserving if, throughout the decision horizon, the graph topology remains exactly unchanged: no nodes or edges are removed, added, rewired, or reweighted. Formally, the graph structure $G=(V,E)$ is invariant over time, and interventions modify only node- or edge-level attributes (e.g., internal opinions, expressed opinions, or fixed-value boundary conditions), while keeping the adjacency structure fixed.
In topology-preserving settings, the observable graph connectivity does not directly reveal which nodes have been intervened on. Consequently, two intervention histories that differ in selected nodes may induce identical or nearly identical adjacency structures. This creates a representation ambiguity under permutation-invariant message-passing encoders, where structurally identical graphs with different intervention histories can produce indistinguishable aggregated embeddings.
Such ambiguity leads to a fundamental challenge:
  • History-induced state aliasing: different intervention trajectories may correspond to states that are structurally identical but behaviorally distinct.
The MI and ME problems defined below are canonical examples of topology-preserving moderation, as they modify internal or expressed opinions without altering the underlying network structure.
3.4 Canonical Moderation under One-Shot Planning with ANP Evaluation
We now formalize the two key intervention problems for reducing polarization and integrate them with the planning constraint and evaluation criterion that define the final setting studied in this work.
Throughout, we assume a budget $k$, and the goal is to select $k$ individuals to moderate, i.e., to set the targeted opinion attribute to a neutral state (0), such that the resulting polarization index is minimized.
@@ 3.4.2 The ModerateExpressed Problem
While the MI and ME problems above specify what is being intervened on, they do not specify how intervention sequences are constructed nor how they are evaluated over time. We now integrate them with the planning and evaluation regime that defines the final decision problem studied in this work.
One-shot planning constraint (plan once from the initial instance).
@@
Accumulated normalized polarization (ANP).
@@
so the full sequence $(v_{1},\dots,v_{k})$ traces a polarization trajectory from $x_{0}=0$ (no intervention) to $x_{k}=1$ (all $k$ interventions executed).
Recall that the polarization index $\pi(\mathbf{z})=\|\mathbf{z}\|_{2}^{2}/n$ is already normalized by the graph size $n$. Following the trajectory-level normalization idea used in accumulated normalized connectivity (ANC) [3], we further define the stepwise normalized polarization score as
$\widetilde{\pi}\!\left(\mathbf{z}^{(t)}\right)=\frac{\pi\!\left(\mathbf{z}^{(t)}\right)}{n}=\frac{\|\mathbf{z}^{(t)}\|_{2}^{2}}{n^{2}}.$ (5)
We then define the accumulated normalized polarization (ANP) score as
$\mathcal{R}_{\mathrm{ANP}}(v_{1},\dots,v_{k})=\frac{1}{k+1}\sum_{t=0}^{k}\widetilde{\pi}\!\left(\mathbf{z}^{(t)}\right).$ (6)
The quantity $\mathcal{R}_{\mathrm{ANP}}$ can be viewed as a discrete approximation to the area under the polarization trajectory curve over the normalized horizon $x\in[0,1]$. Lower values indicate that polarization is reduced earlier and more persistently throughout the intervention trajectory.
@@ 3.5 Computational Complexity
3.6 Extended Tasks Beyond the Canonical FJ Setting
The MI and ME problems defined above correspond to the canonical unweighted setting under the linear Friedkin–Johnsen (FJ) model, with binary internal opinions and topology-preserving interventions. To broaden the moderation problem space while remaining within the same overall framework, we define three extended task families that relax different aspects of this canonical setting. These extensions modify either the opinion representation, the intervention cost structure, or the intervention mechanism, while preserving the objective of minimizing polarization under a fixed budget.
Extended Task I: Cost-Aware Moderation.
In the canonical MI and ME problems, all interventions are assumed to have equal cost. In this extended task, each node $v\in V$ is associated with a nonnegative intervention cost $c(v)$, representing the relative expense or difficulty of moderating that node.
The planner is still given a budget horizon of $k$ interventions, but the quality of a solution now depends not only on the achieved polarization reduction, but also on how effectively the intervention sequence uses the available costs over the trajectory. This extension preserves the underlying FJ opinion dynamics and graph topology, but changes the optimization landscape by introducing heterogeneous action costs.
Extended Task II: Continuous-Opinion Moderation.
In the canonical setting, internal opinions are binary and polarized, i.e., $s_{i}\in\{-1,+1\}$. In this extended task, internal opinions are allowed to take continuous values in the interval [1,1][-1,1].
This extension preserves the same linear FJ opinion dynamics and the same moderation actions as in MI and ME, but relaxes the opinion representation to allow heterogeneous opinion intensities. As a result, users can differ not only in camp membership (negative vs. positive), but also in the strength of their stance. The moderation action still sets the targeted attribute to the neutral value 0, but the resulting intervention problem now depends on a richer continuous opinion landscape.
Extended Task III: Topology-Altering Moderation (Node Removal).
In the canonical MI and ME problems, interventions are topology-preserving: the graph $G=(V,E)$ remains unchanged. In this extended task, an intervention removes a selected node from the network.
Formally, selecting a node $v$ updates the graph as
$G\leftarrow G\setminus\{v\},$
removing $v$ and all its incident edges. Opinion dynamics are subsequently computed on the modified topology.
Unlike MI and ME, this task alters the graph structure itself and therefore changes the influence propagation pattern after each action. The objective remains minimizing polarization under a budget of $k$ interventions, but the feasible action space now includes structural modification of the network.
Unified Problem View.
Under all three extended task families, the core objective remains the same: given a graph, an opinion state, and a budget of $k$ interventions, select an ordered intervention sequence that minimizes polarization.
What changes across tasks is the transition structure of the environment. In cost-aware moderation, the intervention dynamics are unchanged but actions carry heterogeneous costs. In continuous-opinion moderation, the underlying FJ dynamics remain linear, but the opinion space becomes richer by allowing heterogeneous opinion magnitudes. In topology-altering node removal, each action changes the graph itself, modifying future influence propagation and making long-horizon structural effects central to the decision process.
Together, these tasks define an expanded moderation problem space that extends the canonical MI/ME formulation along three complementary dimensions: cost structure, opinion representation, and intervention mechanism. This provides a broader testbed for evaluating whether a moderation framework can generalize beyond the original binary, unweighted, topology-preserving setting while preserving the same overall goal of polarization reduction.
@@ 4 PACIFIER Framework
PACIFIER is a graph-based learning framework for sequential polarization moderation on networks. It is designed to (i) learn intervention policies from simulated graph–opinion interactions and (ii) deploy them efficiently on large real-world graphs via lightweight feed-forward inference. PACIFIER follows a two-phase workflow: (i) offline training on a distribution of synthetic graph–opinion instances, and (ii) deployment-time application on a target network without iterative recomputation during planning.
@@
PACIFIER supports two closely related variants within the same architecture: PACIFIER-RL, which learns a long-horizon intervention policy via bootstrapped value-based reinforcement learning, and PACIFIER-Greedy, a myopic variant that learns to approximate immediate polarization reduction without bootstrapping. Both variants share the same encoder–decoder representation and differ only in their learning objectives and exploration strategies (Sec. 4.4).
@@
Overview and components. Moderation under MI/ME-type interventions typically differs from topology-altering structural control: in many practical settings the graph topology remains fixed, while polarization evolves through opinion dynamics and the history of past interventions. This setting motivates PACIFIER’s design around three core components, each addressing a distinct bottleneck for scalable sequential decision making.
Two-echo-chamber training instances (structured initialization).
Sequential policies require diverse training trajectories, yet purely random graphs and opinion assignments often fail to expose meaningful polarized regimes. PACIFIER therefore trains on a structured distribution of synthetic instances with two opposing communities (echo chambers) and aligned initial opinions (Sec. 4.1). By controlling cross-group exposure (inter-group edge ratio), we generate a spectrum of polarization regimes, which improves robustness and inductive generalization from small training graphs to large test graphs.
Encoder (history-aware and polarization-aware state representation).
In topology-preserving moderation, the adjacency alone cannot reveal which nodes have been intervened on. Consequently, distinct intervention histories can appear nearly identical to permutation-invariant message-passing encoders, leading to history-induced state aliasing and unstable value learning. PACIFIER’s encoder explicitly encodes intervention history through temporal-aware node marking, and augments learned embeddings with polarization-related auxiliary features that summarize opinion–structure interaction patterns (Sec. 4.2). This provides value-relevant global signals without requiring repeated steady-state recomputation as part of the policy input.
Decoder (parallel action scoring for scalable deployment).
At each step, the decision is to select one feasible node to intervene on. To scale to large graphs, PACIFIER uses a value-based decoder that scores all candidate nodes in parallel from shared node/graph embeddings and auxiliary features (Sec. 4.3). A feasibility mask excludes previously intervened nodes, enabling efficient greedy selection at deployment time.
Together, these modules enable PACIFIER to learn intervention strategies offline and execute them online with lightweight feed-forward planning, while remaining compatible with multiple moderation mechanisms and dynamics (Sec. 4.5).